A runner travels 1.5 laps around a circular track in a time of 50 s. The diameter of truck is 40 m and its circumference is 126 m. Find (a) the average speed of the runner and (b) the magnitude of the runner averwge's velocity. Be careful here: average speed depends on the distance traveled, whereas average depends on the displacement at the end of the particular journey.


1
Expert's answer
2021-03-10T17:22:00-0500

a) The distance travelled by the runner (1.5 of the circle with the circumference 126 m) is:


"l = 1.5\times 126m = 189m"

Since the time is "t = 50s", then the average speed is:


"speed = \dfrac{l}{t} = \dfrac{189m}{50s} = 3.78 m/s"

b) The displacement of the runner at the end of the jorney is equal to the diameter of the circle, since she completed 1.5 laps. The diameter is "d = 40m". Thus, the average velocity is:


"velocity = \dfrac{d}{t} = \dfrac{40m}{50s} = 0.8m/s"

Answer. a) 3.78 m/s, b) 0.8 m/s.